OpenStudy (anonymous):

The value of the dependent variable in a function is always ______ the value of independent variable A. determined by B. greater than C.equal to D. less than

OpenStudy (anonymous):

please help

OpenStudy (freckles):

let's think about it y=f(x) f(x)=x^2 here are few points on that graph: (x,y) ---- (1/2,1/4) (0,0) (2,4) is the x value always equal to the y value? is the x value always greater than the the y value? is the x value always less than the the y value? does the value I pick for x determine the value of y?

OpenStudy (anonymous):

so d answer is

OpenStudy (freckles):

so you think y is less than x always? well 1/4 is less than 1/2 but is 0 less than 0 or is 4 less than 2?

OpenStudy (freckles):

do you think 4 is less than 2?

OpenStudy (anonymous):

no

OpenStudy (freckles):

so that is a counterexample for your choice y being less than x always

OpenStudy (freckles):

so the points I have you see that x isn't always equal to y since 1/2 certainly doesn't equal 1/4 or 2 doesn't equal 4. also y isn't always greater than x in my points I have above since in the point (1/2,1/4) 1/4 isn't greater than 1/2 and also the other point we have (0,0) 0 isn't greater than 0

OpenStudy (freckles):

so does the value of x we pick determine the value of y? well I picked all those x values above to determine what y would be for each of those x's

OpenStudy (freckles):

anyways hope that helps i must leave you now

OpenStudy (anonymous):

k

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so think about the words. the DEPENDENT variable and the INDEPENDENT variable. Your grades are determined by the amount of effort you put forth. so they are dependent on your effort. effort would be independent and grades would be DEPENDENT since they rely on effort, right?

OpenStudy (anonymous):

but i think its c

OpenStudy (anonymous):

not always

OpenStudy (freckles):

@jcooper344 I guess you still don't understand what I said above.... do you know in f(x)=x^2 (f(x)=y) that x is the independent variable and y is the dependent variable (x,y) is an ordered pair showing the y value that corresponds to the x value we pick I selected through x's to plug into my function f(x)=x^2 f(1/2)=(1/2)^2=(1/2)(1/2)=1/4 f(0)=0^2=0(0)=0 f(2)=2^2=2(2)=4 so the three ordered pairs I wanted you to look at there was: (1/2,1/4) (0,0) (2,4) the first number in each of those pairs is x (the independent variable) the second number in each of those pairs is y (the dependent variable) then I asked you the following (look at the pairs by the way): is y always greater than x? (hint: look at the 1st and 2nd pairs) is y always smaller than x? (hint: look at the 2nd and 3rd pairs) is y always equal to x? (hint: look at the 1st and 3rd pairs) does y depend on x? (hint: notice how I found the y values for each of the 3 x's I chose)

OpenStudy (freckles):

also here a major hint in the words dependent think depends on

OpenStudy (anonymous):

nope but im done